Grade 12Algebra

show that the points A(1,0,-6) ,B(-3,6,2) and C(-5,9,6) are collinear

ankur
8 Years agoGrade 12

1 Answer

Khimraj
8 Years ago
First find line passing through point A and B. If C satisfy it then points are collinear.
A(1,0,-6) and B(-3,6,2)
direction vector (4,-6,-8)
equation of line AB
(x-1)/4  = (y-0)/(-6)  =  (z+6)/(-8)
It satisfies point C.
so point A,B,C are collinear.
Hope it clears. If you like answer then please approve it.